OpenStudy (anonymous):

A constant force if 600 N is applied straight up on a 500 kg crate. How long will it take for the crate to rise 8.0 m from rest?

OpenStudy (anonymous):

as F=MA i.e. mass*acceleration so from the given data A=F/M=600/500=1.2ms^-2 using equation of motions i.e. s=ut+1/2(at^2) where s= displacement in meter u=initial speed t=time a=acceleration so, 8=1/2(1.2*t^2)>>>>as u=0 given at rest initially 16/1.2=t^2 t=3.65148seconds :)

OpenStudy (anonymous):

can you explain the first equation? i'm kinda having a brain fart! D:

OpenStudy (anonymous):

newton law F=ma force=mass*acceleration